A line that is perpendicular to 3x-9y=13 through (2,-4)this is what I did:
-9y=13-3x
divided all by -9 to get
y=-13/9-1/3x
slope =-1/3
y-(-4)=-1/3(x-2)
y+4=-1/3x+2/3
y=-1/3x+2/3-4
y=-1/3x-3 1/3

Answer by checkley71(8403)

SORRY FOR THE OVERSIGHT. I JUST HAD LASER SURGERY THIS MORNING & MY LEFT EYE IS STILL BLURRED. HOWEVER I SHOULD HAVE RECHECKED MY MATH.
3X-9Y=13 OR -9Y=3X-13 OR -Y=3/9X-13/9 OR -Y=1/3X-13/9 or Y=-1/3X+13/9 SO THE PERPENDICULAR LINE HAS A SLOPE OF 3 THE NEGATIVE RECIPRICAL OF THE SLOPE OF
-1/3
SO THE EQUATION OF THE LINE WITH A SLOPE OF 3 THROUGH (2,-4) IS
Y=3X+b or -4=3*2+b or -4=6+b or b=-10 thus the equatiion is y=3x-10